# Enhancing Language Model Factuality via Activation-Based Confidence Calibration and Guided Decoding

### v1

@@ Abstract
Calibrating language models (LMs) aligns their generation confidence with the actual likelihood of answer correctness, which can inform users about LMs’ reliability and mitigate hallucinated content. However, prior calibration methods, such as self-consistency-based and logit-based approaches, are either limited in inference-time efficiency or fall short of providing informative signals. Moreover, simply filtering out low-confidence responses reduces the LM’s helpfulness when the answers are correct. Therefore, effectively using calibration techniques to enhance an LM’s factuality remains an unsolved challenge. In this paper, we first propose an activation-based calibration method, ActCab, which trains a linear layer on top of the LM’s last-layer activations that can better capture the representations of knowledge. Built on top of ActCab, we further propose CoDec, a confidence-guided decoding strategy to elicit truthful answers with high confidence from LMs. By evaluating on five popular QA benchmarks, ActCab achieves superior calibration performance than all competitive baselines, e.g., by reducing the average expected calibration error (ECE) score by up to 39%. Further experiments on CoDec show consistent improvements in several LMs’ factuality on challenging QA datasets, such as TruthfulQA, highlighting the value of confidence signals in enhancing the factuality.
@@ 1 Introduction
Using ActCab, our second contribution is a Confidence-guided Decoding strategy, CoDec. CoDec directs the LM to generate outputs with higher confidence by incorporating the confidence scores of candidate tokens with the top-K highest probabilities predicted by the LM. This increases the likelihood that the answers are correct while maintaining inference efficiency. Unlike selective generation, CoDec preserves the helpfulness of LM by not filtering out responses, but rather by better discerning the correct answers. This aligns with the principles of other inference-time intervention methods (Li et al., 2023; Zou et al., 2023; Bayat et al., 2024). However, CoDec differs by keeping the model unchanged and does not affect the LM’s original reasoning process, resulting in more stable performance across different training data and LMs, as shown in our experiments.
@@ 2.2 Enhancing Factuality via Inference Intervention
Inference intervention methods collect directional vectors representing truthfulness and incorporate them into the LLMs’ forward pass, thereby steering them towards factual outputs. For instance, Inference-time Intervention (ITI) (Li et al., 2023) uses linear probing to identify attention heads that exhibit distinct activation distributions for true and false statements. Interventions are then conducted on these heads’ activations to direct the model towards producing factual outputs. Similarly, Representation Engineering (RepE) (Zou et al., 2023) detects truthful directions in each layer by prompting the language model with pairs of instructions that have contrasting meanings and incorporates these directions into each layer during decoding. These methods are closely related to CoDec, since they also modify model behavior during the decoding phase. The distinction between CoDec and intervention methods lies in their application. Both ITI and RepE directly intervene in the activations of LMs, while CoDec adjusts the LM’s output distribution. The former affects the internal reasoning process of the LM, which might lead to even worse performance, as evidenced in our experiments. Conversely, CoDec does not affect the model’s inherent reasoning process, keeping the LM unchanged and providing stable improvement.
@@ 3.1 Eliciting Confidence from Activations
$\mathcal{L}_{MSE}=\frac{1}{|\mathcal{D}|}{\textstyle\sum_{(x,y)\in\mathcal{D}}}(\mathbf{1}(y\text{ is correct})-p_{\theta}(y|x))^{2},$ (1)
@@ 3.2 Constructing Soft Training Labels via $K$-fold Cross-validation
$\mathcal{L}_{ECE}=\frac{1}{|\mathcal{D}|}{\textstyle\sum_{(x,y)\in\mathcal{D}}}(\operatorname{acc}\left(B_{(x,y)}\right)-p_{\theta}(y|x))^{2}$ (3)
@@ 4 CoDec: Confidence-guided Decoding
Intuitively, a response generated with higher confidence is more likely to be correct, assuming the LM is well-calibrated. Building on this intuition, we design CoDec, a decoding strategy that guides the LM to generate high-confidence responses. We illustrate the whole process of CoDec in Figure 2.
@@ 5.1 Datasets
We follow Liu et al. (2024) to assess the calibration performance of ActCab and baseline methods using CaT, a public calibration evaluation benchmark. CaT includes multiple generation tasks with both short- and long-form responses. In our experiments, we focus on the phrase- and sentence-level tasks in CaT, in particular, TruthfulQA (Lin et al., 2022b), TriviaQA (Joshi et al., 2017), SciQ (Johannes Welbl, 2017), Natrual Questions (NQ) (Kwiatkowski et al., 2019), and WikiQA (Yang et al., 2015). The statistics for each dataset are presented in Table 1. Additionally, we evaluate the factuality of LMs on these five tasks.
@@ 5.2 Implementation Details
We implement the in-context learning setting in all our experiments. Following Liu et al. (2024), we select 15 QA pairs from the training set as demonstrations when evaluating the calibration methods. To assess the factuality of CoDec and the baseline methods, we use a 5-shot generation setting, consistent with the in-context learning setting used in ITI (Li et al., 2023). To find the optimal hyperparameters for ActCab, CoDec, and comparisons, we use 20% of the training samples as a validation set and train the classifier on the remaining instances. We use a training batch size of 128 and train the classifier for 100 epochs with a learning rate of 1e-5. We set $\lambda$ in Equation 4 to 0.3. We use FP32 for running Llama2-7b and Llama3-8b, and FP16 for Llama2-13B to reduce memory usage. Since ActCab only trains a single linear layer, the training process takes less than 1 hour on a single A6000 GPU for each task. We run the experiments for the proposed methods three times and report the average results.
@@
Table 2: Results of ActCab and the baselines on CaT, with the best results in bold. The last row shows the averaged metric values across five tasks. ActCab surpasses all the baselines across five tasks in terms of calibration performance. The ablation study (ActCab w/o ECE loss, last column) demonstrates the effectiveness of ECE loss. We highlight numbers where ActCab improves over both the Seq. Likelihood and all baselines in blue; when ActCab outperforms the Seq. Likelihood, it is colored in green.
@@ 5.3 Evaluation Metrics
In addition to ECE, we use Brier Score to measure the calibration performance of ActCab and baselines. Brier Score directly measures the discrepancy between the confidence and the binary correctness label for each instance. It is computed as $Brier=\frac{1}{N}{\textstyle\sum_{y\in Y}(f(x,y)-\mathbf{1}(y\text{ is correct}))}$.
@@ 5.4 Baselines
Table 3: Factuality results of CoDec and baselines on five tasks, with the best results in bold. Greedy Decoding refers to the plain LM with in-context learning prompting using greedy decoding. We highlight numbers where ActCab improves over both the Greedy Decoding and all baselines in blue. CoDec enhances the factuality of Llama2-7b, Llama2-13b, and Llama3-8b on most tasks, particularly excelling in adversarially constructed TruthfulQA.
@@ 5.5.2 Results on LM Factuality
However, CoDec does not achieve satisfactory performance on SciQ on Llama2-7b and Llama3-8b. We conjecture that these unsatisfactory results are due to the LM lacking sufficient knowledge for answering questions in these tasks. Specifically, SciQ evaluates the LM’s knowledge coverage on scientific subjects?. However, the classifier that CoDec uses, a single linear layer, cannot capture sufficient knowledge to answer these scientific questions. In contrast, the adversarially constructed TruthfulQA challenges the LMs with popular misconceptions, where the models have seen the knowledge during pretraining and CoDec plays a role of steering the LM to output the correct facts. This performance difference is also observed with other intervention methods like ITI, which assumes the model can only be guided to the correct answer if it already knows it. To illustrate these points, we provide two examples in Table 4. As can be seen, CoDec does not help the LM answer the scientific question “There are about 6200 known species of what?”, as it might lack this specific knowledge. However, CoDec prevents the LM from propagating misconceptions like “Einstein flunked math and physics.”
@@
Table 4: Examples of greedy decoding and CoDec response in SciQ and TruthfulQA. Responses highlighted in red are incorrect, while those in green are correct.
@@ 6 Conclusion
This work presents methods aimed at improving factuality and reducing harmful responses in LM outputs through Activation-Based Confidence Calibration (ActCab) and Confidence-Guided Decoding (CoDec). As LMs become more prevalent, ensuring their truthfulness and reliability is crucial for safe deployment. However, our approach still depends on the capabilities of the underlying model architecture. Future research must continue to address potential harms, such as bias, toxicity, and misinformation, as well as the risks of adversarial misuse and privacy concerns.

### v2

@@ Abstract
Calibrating language models (LMs) aligns their generation confidence with the actual likelihood of answer correctness, which can inform users about LMs’ reliability and mitigate hallucinated content. However, prior calibration methods, such as self-consistency-based and logit-based approaches, are either limited in inference-time efficiency or fall short of providing informative signals. Moreover, simply filtering out low-confidence responses reduces the LM’s helpfulness when the answers are correct. Therefore, effectively using calibration techniques to enhance an LM’s factuality remains an unsolved challenge. In this paper, we first propose an activation-based calibration method, ActCab, which trains a linear layer on top of the LM’s last-layer activations that can better capture the representations of knowledge. Built on top of ActCab, we further propose CoDec, a confidence-guided decoding strategy to elicit truthful answers with high confidence from LMs. By evaluating on five popular QA benchmarks, ActCab achieves superior calibration performance than all competitive baselines, e.g., by reducing the average expected calibration error (ECE) score by up to 39%. Further experiments on CoDec show consistent improvements in several LMs’ factuality on challenging QA datasets, such as TruthfulQA, highlighting the value of confidence signals in enhancing the factuality.  [^1] Code is available at https://github.com/launchnlp/ActCab.
@@ 1 Introduction
Using ActCab, our second contribution is a Confidence-guided Decoding strategy, CoDec. CoDec directs the LM to generate outputs with higher confidence by incorporating the confidence scores of candidate tokens with the top-K highest probabilities predicted by the LM. This increases the likelihood that the answers are correct while maintaining inference efficiency. Unlike selective generation, CoDec preserves the helpfulness of LM by not filtering out responses, but rather by better discerning the correct answers. This aligns with the principles of other inference-time intervention methods (Li et al., 2023; Zou et al., 2023; Fatahi Bayat et al., 2024). However, CoDec differs by keeping the model unchanged and does not affect the LM’s original reasoning process, resulting in more stable performance across different training data and LMs, as shown in our experiments.
@@ 2.2 Enhancing Factuality via Inference Intervention
Inference intervention methods collect directional vectors representing truthfulness and incorporate them into the LMs’ forward pass, thereby steering them towards factual outputs. For instance, Inference-time Intervention (ITI) (Li et al., 2023) uses linear probing to identify attention heads that exhibit distinct activation distributions for true and false statements. Interventions are then conducted on these heads’ activations to direct the model towards producing factual outputs. Similarly, Representation Engineering (RepE) (Zou et al., 2023) detects truthful directions in each layer by prompting the language model with pairs of instructions that have contrasting meanings and incorporates these directions into each layer during decoding. These methods are closely related to CoDec, since they also modify model behavior during the decoding phase. The distinction between CoDec and intervention methods lies in their application. Both ITI and RepE directly intervene in the activations of LMs, while CoDec adjusts the LM’s output distribution. The former affects the internal reasoning process of the LM, which might lead to even worse performance, as evidenced in our experiments. Conversely, CoDec does not affect the model’s inherent reasoning process, keeping the LM unchanged and providing stable improvement.
@@ 3.1 Eliciting Confidence from Activations
$\mathcal{L}_{MSE}=\frac{1}{|\mathcal{D}|}{\sum_{(x,y)\in\mathcal{D}}}(\mathbf{1}(y\text{ is correct})-p_{\theta}(y|x))^{2},$ (1)
@@ 3.2 Constructing Soft Training Labels via $K$-fold Cross-validation
$\mathcal{L}_{ECE}=\frac{1}{|\mathcal{D}|}{\sum_{(x,y)\in\mathcal{D}}}(\operatorname{acc}\left(B_{(x,y)}\right)-p_{\theta}(y|x))^{2}$ (3)
@@ 4 CoDec: Confidence-guided Decoding
Intuitively, a response generated with higher confidence is more likely to be correct, assuming the LM is well-calibrated. Building on this intuition, we design CoDec, a decoding strategy that guides the LM to generate high-confidence responses, in line with recent guided decoding studies (Khalifa et al., 2023). We illustrate the whole process of CoDec in Figure 2.
@@ 5.1 Datasets
We assess the calibration performance of ActCab and baseline methods using CaT, a public calibration evaluation benchmark (Liu et al., 2024). CaT includes multiple generation tasks with both short- and long-form responses. In our experiments, we focus on the phrase- and sentence-level tasks in CaT, in particular, TruthfulQA (Lin et al., 2022b), TriviaQA (Joshi et al., 2017), SciQ (Johannes Welbl, 2017), Natrual Questions (NQ) (Kwiatkowski et al., 2019), and WikiQA (Yang et al., 2015). The statistics for each dataset are presented in Table 1. Additionally, we evaluate the factuality of LMs on these five tasks.
@@ 5.2 Implementation Details
We implement the in-context learning setting in all our experiments. Following Liu et al. (2024), we select 15 QA pairs from the training set as demonstrations when evaluating the calibration methods. To assess the factuality of CoDec and the baseline methods, we use a 5-shot generation setting, consistent with the in-context learning setting used in ITI (Li et al., 2023). To find the optimal hyperparameters for ActCab, CoDec, and all comparisons, we follow Liu et al. (2024) to use 20% of the training samples as a validation set and train the classifier on the remaining instances. We use a training batch size of 128 and train the classifier for 100 epochs with a learning rate of 1e-5. We set $\lambda$ in Equation 4 to 0.3. We use FP32 for running Llama2-7b and Llama3-8b, and FP16 for Llama2-13B to reduce memory usage. Since ActCab only trains a single linear layer, the training process takes less than 1 hour on a single A6000 GPU for each task. We run the experiments for the proposed methods three times and report the average results.
@@
Table 2: Results of ActCab and comparison methods on CaT, with the best results in bold. The last row shows the averaged metric values across five tasks. ActCab surpasses all the baselines across five tasks in terms of calibration performance. The ablation study (ActCab w/o ECE loss, last column) demonstrates the effectiveness of ECE loss. We highlight numbers where ActCab improves over both the Seq. Likelihood and all baselines in blue; when ActCab outperforms the Seq. Likelihood, it is colored in green.
@@ 5.3 Evaluation Metrics
In addition to ECE, we use Brier Score to measure the calibration performance of ActCab and baselines. Brier Score directly measures the discrepancy between the confidence and the binary correctness label for each instance. It is computed as $Brier=\frac{1}{N}{\textstyle\sum_{y\in Y}(f(x,y)-\mathbf{1}(y\text{ is correct}))^{2}}$.
@@ 5.4 Baselines
Table 3: Factuality results of CoDec and comparisons on five tasks, with the best results in bold. Greedy Decoding refers to the plain LM with in-context learning prompting using greedy decoding. We highlight numbers where CoDec improves over both the Greedy Decoding and all baselines in blue. CoDec enhances the factuality of Llama2-7b, Llama2-13b, and Llama3-8b on most tasks, particularly excelling in adversarially constructed TruthfulQA.
@@ 5.5.2 Results on LM Factuality
However, CoDec does not achieve satisfactory performance on SciQ on Llama2-7b and Llama3-8b. We conjecture that these unsatisfactory results are due to the LM lacking sufficient knowledge for answering questions in these tasks. Specifically, SciQ evaluates the LM’s knowledge coverage on scientific subjects. However, the classifier that CoDec uses, a single linear layer, cannot capture sufficient knowledge to answer these scientific questions. In contrast, the adversarially constructed TruthfulQA challenges the LMs with popular misconceptions, where the models have seen the knowledge during pretraining and CoDec plays a role of steering the LM to output the correct facts. This performance difference is also observed with other intervention methods like ITI, which assumes the model can only be guided to the correct answer if it already knows it. To illustrate these points, we provide two examples in Table 4. As can be seen, CoDec does not help the LM answer the scientific question “There are about 6200 known species of what?” as it might lack this specific knowledge. However, CoDec prevents the LM from propagating misconceptions like “Einstein flunked math and physics.”
@@
Table 4: Examples of greedy decoding and CoDec response in SciQ and TruthfulQA. Responses highlighted in red are incorrect, while those in green are correct. More failed examples of CoDec in SciQ are provided in Appendix C.
@@ 5.5.3 Robustness across Training Data Sources and LMs
Deocding Strategy Greedy Search ITI CoDec
Throughputs (Toks/s) 29.07 28.79 24.76
True.*Info. 27.50 27.00 41.10
Table 6: Throughputs of decoding methods on TruthfulQA using a single A100 GPU, measured in tokens generated per second. The base LM is Llama2-7b. CoDec improves True.*Info. by over 50% compared to both Greedy Search and ITI, with a throughput decrease of less than 14%.
5.6 Decoding Speed
Compared to the greedy search baseline, CoDec introduces additional computation by obtaining the hidden states of candidate tokens. To reduce decoding latency, we apply two strategies to speed up CoDec: (i) we batch the 7 candidate tokens when retrieving their hidden states, enabling parallel processing; (ii) we reuse the hidden state of the selected candidate token from the previous step when computing the token probabilities at the current step, avoiding the need to recompute hidden states by passing through the entire model again. For quantitative analysis, we compare the throughputs of CoDec with greedy search and ITI on TruthfulQA, and list the results in Table 6.333Experiments are conducted using Huggingface Transformers https://github.com/huggingface/transformers. We can observe that CoDec results in less than a 14% reduction in throughput compared to ITI, while improving True.*Info by over 50%. This makes CoDec particularly suitable for high-stakes domains, where accuracy is critical and latency is a secondary concern, such as finance and legal.
@@ 6 Conclusion
Acknowledgements
This work is supported in part by National Science Foundation through grant 2046016. We are grateful to all ARR reviewers for providing useful feedback.
@@ Acknowledgements
This work presents methods aimed at improving factuality and reducing harmful responses in LM outputs through activation-based confidence calibration (ActCab) and confidence-guided decoding (CoDec). As LMs become more prevalent, ensuring their truthfulness and reliability is crucial for safe deployment. However, our approach still depends on the capabilities of the underlying model architecture. Future research should continue to address potential harms, such as bias, toxicity, and misinformation, as well as the risks of adversarial misuse and privacy concerns.
@@ Appendix B Prompt for TruthfulQA
Task Example
SciQ
Question: About how tall can mid-ocean ridges be?
Ground Truth: about 2 km
Greedy Decoding: 10000 meters
CoDec: 1000 meters
Question: What is the term for the total kinetic energy of moving particles of matter?
Ground Truth: thermal energy
Greedy Decoding: kinetic energy
CoDec: temperature
Question: What do most living things use to make atp from glucose?
Ground Truth: oxygen
Greedy Decoding: ATP synthase
CoDec: ATP synthase
Question: What is the most abundant metal of the earth’s crust?
Ground Truth: aluminum
Greedy Decoding: iron
CoDec: iron
Table 7: Failure examples of CoDec in SciQ.
Appendix C Failed examples of CoDec response in SciQ
Table 7 presents several failure examples of CoDec. These cases echo our hypothesis that CoDec struggles due to the language model’s lack of scientific knowledge. As a result, CoDec, which consists of an additional single linear layer, either generates an incorrect answer (as seen in the first two examples) or repeats the same incorrect response as the greedy search (as in the third and fourth examples).